An airplane has a velocity relative to the air of 223 m/s in a westerly direction. If the wind has a speed relative to the ground of 69 m/s directed to the north, what is the speed of the airplane relative to the ground?
Answer #1

SPEED OF THE PLANE RELATIVELY TO THE GROUND = PALNE VELOCITY WITH RESPECT TO THE AIR+ VELOCITY OF AIR WITH RESPECT TO THE GROUND

Vp/g = Vp/a+Va/g = -223i-69j

magnitude = sqrt(223^2+69^2) = 233.43 m/sec

direction = tan^-1(69/223) = 17.19 degrees